# CommVQA: Situating Visual Question Answering in Communicative Contexts

### v1

@@ 5 Model Benchmarking
Model BLEU-1 BLEU-4 METEOR ROUGE CIDEr CLIPScore
IDEFICS 0.326 0.106 0.189 0.401 0.825 0.667
IDEFICS (Contextual) 0.356 0.122 0.198 0.413 0.904 0.666
mPLUG-Owl 0.191 0.045 0.166 0.311 0.414 0.720
mPLUG-Owl (Contextual) 0.178 0.043 0.169 0.284 0.335 0.756
LLaVA 0.192 0.050 0.171 0.311 0.431 0.729
LLaVA (Contextual) 0.149 0.037 0.161 0.247 0.200 0.784
BLIP-2 0.289 0.067 0.100 0.286 0.435 0.604
BLIP-2 (Contextual) 0.059 0.010 0.078 0.224 0.343 0.662
Table 1: Results for the baseline and contextual condition across all models. The highest-scoring models in all columns are the ones which integrate the contextual condition. The CLIPScore metric improves for all models, except for IDEFICS, where it stays the same. This indicates the model’s generated answers contain more visual details with the contextual condition. Results were conducted over three confidence intervals.
In this section, we investigate the performance of four state-of-the-art VLMs on CommVQA  and to what extent providing context to the models improves their performance.
We selected four models to benchmark on this dataset: LLaVA Liu et al. (2023), BLIP-2 Li et al. (2023), mPLUG-OWL Ye et al. (2023), and IDEFICS Laurençon et al. (2023). To maximally enable the reproducibility of this work and support the public development of models, we focus on open-source models. LLaVA is trained on a multimodal instruction-following dataset, and exhibits strong performance across several VQA datasets Goyal et al. (2017); Hudson and Manning (2019). mPLUG-Owl Ye et al. (2023) is also instruction-tuned, and displays competitive performance on various VQA benchmarks, including VQA-v2 Goyal et al. (2017). IDEFICS Laurençon et al. (2023) is an open-access reproduction of Flamingo Alayrac et al. (2022), and was trained on a naturalistic web-scale dataset of interleaved image-text documents, including 141 million web pages. We selected LLaVA, mPLUG-Owl, and IDEFICS since their instruction-tuned nature allows us to straightforwardly integrate context. We also included a non-instruction-tuned model, BLIP-2 Li et al. (2023), a general-purpose vision-language model that is additionally fine-tuned on the VQA task.
@@
We evaluated these models on the NLG metrics BLEU Papineni et al. (2002), CIDEr Vedantam et al. (2015), METEOR Banerjee and Lavie (2005), and ROUGE Lin (2004). However, there are still many aspects of performance these metrics cannot capture. For instance, studies show that these metrics do not adequately capture object hallucinations Rohrbach et al. (2018). To measure how closely an answer’s visual details matches with the image, we also include the CLIPScore metric Hessel et al. (2021). Including the CLIPScore metric has the added benefit of including an assessment of how close a model-generated answer is to a visual description.
@@
For all models, we ran baseline experiments where only the image and question were provided, and the full contextual condition where the scenario and description were additionally provided, just like for the participants who answered the questions. We include our prompt for the contextual condition in Appendix A.1. All evaluations were conducted with the greedy decoding method to ensure reproducibility.
@@
In Table 1, we show the performance of all models when provided with the image and question (baseline) and the full context that was available to human answerers (contextual).
The IDEFICS model with the contextual condition has the highest performance for all models across all metrics except CLIPScore, which remains unaffected. Evidently, integrating contextual information made the content itself more aligned with the ground-truth answer references, as evaluated by the reference-based metrics. The unchanged CLIPScore rating suggests that the contextually induced changes in the generated answers don’t change how they relate to the images.
While for IDEFICS (the best-performing model), CLIPScore is the only metric that doesn’t significantly improve in the contextual condition, this pattern is reversed for all other models. Here, the contextual integration primarily boosts CLIPScore (i.e., the similarity to the image) but results in worse performance according to the reference-based metrics. One potential explanation for this pattern is that the less well-performing models closely reiterate the visual information they receive from the description in the contextual condition, instead of adding new information to answer the question. This would result in higher CLIPScores but lower similarity with the ground-truth answers.
To investigate how much visual information models re-iterate within their generated answers, we compared the average cosine similarity (using Sentence-BERT Reimers and Gurevych (2019) embeddings) between the description and the model’s generated answers. Figure 4 shows that the similarity between the description and the generated answers significantly increases with the contextual condition across models, according to a two-sample t-test analysis. However, the increase in description similarity between the baseline and contextual condition is the lowest for IDEFICS. We conclude that in their answers, the contextual versions of LLaVA, mPLUG-Owl, and BLIP-2 mimic the description more than IDEFICS does, which explains the CLIPScore increase across the baseline and contextual conditions for these three models. These findings suggest that those models might overly emphasize the descriptive details when available, leading to less alignment with the ground-truth answers.
@@ 5.1 Quantifying Hallucinations
While IDEFICS (contextual) best approximates the human answers, it’s still far from achieving human-like performance. When a model is intended to make information available that can’t be verified by a user, it is especially important that this model only generates truthful content MacLeod et al. (2017). While this is straightforward to evaluate for accuracy-based VQA datasets, this isn’t easily captured by the similarity-based metrics that are used for long-form evaluation, as presented here. In this section, we aim to give a brief intuition about the rate of hallucinated or incorrect pieces of information contained in the answers of the best-performing model, IDEFICS (contextual).
@@
To estimate how many answers contain wrong, hallucinated or unverifiable content generated by the model, three authors independently annotated a randomly sampled subset of 51 generated answers, with a Fleiss inter-annotator kappa agreement of 49.4%. The annotations show that 38.2% of the model-generated answers contained clearly erroneous information and for another 15%, the truthfulness couldn’t be established. These results indicate that even the best-performing model (IDEFICS (contextual)) is still far from human-like performance on CommVQA.
Since named entities are a likely source of error and hallucination for VLMs, we further analyze the potential accuracy of named entity recognition within the dataset at scale. Without the context, only 35.4% of the unique named entities in IDEFICS’ answer also appeared in any of the reference answers, suggesting that IDEFICS is frequently naming incorrect entities. Providing IDEFICS with the context guides the model to identify the named entities in the answers more accurately, at a rate of 51%. Most of the named entities that IDEFICS (contextual) struggles on are numbers, particularly in counting. Of the misidentified named entities, 38% of the misidentified named entities were cardinals, ordinals, and quantity estimations, and 25% were time and date indicators, indicating that this model still needs to improve in counting and quantity estimations.
@@
Figure 4: Average cosine similarity to each description from the human-elicited answers and model-generated answers across all conditions. Asterisks indicate significance levels according to a two-sample t-test analysis.
@@ 6 Conclusion
Acknowledgements
This research is supported in part by grants from Google and the Generative AI for the Future of Learning program at Stanford. We thank our experiment participants for their invaluable input. We also thank Bay Foley-Cox for his thorough feedback on previous drafts, and all members of the Coalas Lab.
@@ Appendix A Appendix
A.1 Dataset Collection Overview
@@
For the context matching stage, we prompted GPT-4V with this prompt:
@@
For the answer elicitation study, we asked participants to: “Imagine you are browsing a {scenario} website when you encounter the following image. {image} Another user cannot see the image directly but has access to the following image description: {description}. Based on the description, they asked a follow-up question. Please answer based on the image. {question}.”
@@
A.2 Additional Dataset Composition Analysis
Figure 5: Overall named entity recognition analysis on answers written by participants. Error bars represent 95% confidence intervals.
A.2.1 Answer analysis of named entities
As reported for the analysis of the questions (Section 4.2), we find that there is also significant variation in the number of named entities referenced in the answers. As Figure 5 illustrates, most organizations and people are referenced in the news context, and the least are mentioned in the health and social media scenarios, highlighting how the contextual variance in the questions trickles down to the answer distributions.
Figure 6: Example entries in the CommVQA dataset.
A.2.2 Answer analysis of polar questions
We find that participants readily volunteer information that are not strictly necessary for answering the question directly. For this purpose, polar questions in VQA are an interesting case study, since for polar questions, only an answer of “Yes” or “No” is strictly required. However, in CommVQA, the average length of answers to polar questions was 52.1 characters. A direct string matching analysis revealed that only 0.01% of the answers to polar questions were simply “Yes” or “No,” indicating that most of the time participants provided further information. For example, the question “Is this something you think I could do with limited sight?” received the answer “You could possibly do this with limited sight, but I wouldn’t suggest it.” Prior work suggests that such strictly overinformative behavior is driven by considerations of what information is relevant to a questioner’s goals Tsvilodub et al. (2023). Consequently, the fact that we’re seeing this behavior could again be attributed to the communicative setup of the task, and suggests that positioning VQA as such results in answer patterns that explicitly consider relevancy to a questioner’s goals.
However, an answer longer than “Yes” or “No” might be only restating what was in the question–e.g., if the question was “Are there any trees in the image?”, an answer might be, “Yes, there are trees in the image.” As an exploratory analysis, we prompted GPT-4 OpenAI (2023a) to classify whether the provided answer strictly adhered to the scope of the question, or whether it went beyond and offered additional details. Surprisingly, we found that out of the 528 polar questions in our dataset, only 34% stopped at providing enough information to answer the question, while the other 63% answers were classified as providing information beyond what was necessary to answer the question.
Of the answers to polar questions that over-informed, after we conducted further GPT-4 OpenAI (2023a) analysis, approximately 39% were describing visual details not directly asked about in the question (e.g., Question: Are there any other objects in the image? Answer: Small fish) while 60% were expressing uncertainty and making an educated guess or inference (e.g., Question: Is the athlete a professional? Answer: It’s not explicitly clear whether or not the athlete is a professional although one can infer that the athlete is an enthusiast who clearly trains for the sport and enjoys participating in competitions).
A.3 Models Chosen and Parameters
For running model evaluation, we used the following model versions:
  1. 1.
    LlaVA: llava-v1.5-13b-3GB
  2. 2.
    IDEFICS: idefics-9b-instruct
  3. 3.
    mPLUG-Owl: mplug-owl-llama-7b.
  4. 4.
    BLIP-2: blip2-opt-2.7b.

### v2

@@ 5 Model Benchmarking
Model BLEU-1 BLEU-4 METEOR ROUGE CIDEr
IDEFICS 0.273 0.084 0.179 0.378 0.758
IDEFICS (Contextual) 0.285 0.092 0.195 0.392 0.839
LLaVA 0.195 0.052 0.172 0.317 0.453
LLaVA (Contextual) 0.152 0.037 0.161 0.253 0.219
mPLUG-Owl 0.199 0.048 0.169 0.320 0.462
mPLUG-Owl (Contextual) 0.187 0.046 0.172 0.295 0.355
BLIP-2 0.267 0.059 0.098 0.282 0.434
BLIP-2 (Contextual) 0.015 0.001 0.028 0.043 0.014
Table 1: Comparison of Baseline and Contextual Conditions Across Models. This table presents results for both baseline and contextual conditions across all models. IDEFICS (contextual) achieved the highest scores across all metrics. Results are averaged over three random data splits and model-generated answers.
In this section, we investigate the performance of four state-of-the-art vision-language models on CommVQA  and to what extent providing context to the models improves their performance.
We selected four models to benchmark on this dataset: LLaVA Liu et al. (2024), BLIP-2 Li et al. (2023), mPLUG-OWL Ye et al. (2023), and IDEFICS Laurençon et al. (2023). To maximally enable the reproducibility of this work and support the public development of models, we focus on open-source models. LLaVA is trained on a multimodal instruction-following dataset, and exhibits strong performance across several VQA datasets Goyal et al. (2017); Hudson and Manning (2019). mPLUG-Owl Ye et al. (2023) is also instruction-tuned, and displays competitive performance on various VQA benchmarks, including VQA-v2 Goyal et al. (2017). IDEFICS Laurençon et al. (2023) is an open-access reproduction of Flamingo Alayrac et al. (2022), and was trained on a naturalistic web-scale dataset of interleaved image-text documents, including 141 million web pages. We selected LLaVA, mPLUG-Owl, and IDEFICS since their instruction-tuned nature allows us to straightforwardly integrate context. We also included a non-instruction-tuned model, BLIP-2, a general-purpose vision-language model that is additionally fine-tuned on the VQA task.
Figure 3: Example of Context Dependency in Answer Generation. In this example, the question explicitly asks for content that is not in the description. While the human-elicited answers do not repeat information in the description, IDEFICS (contextual) provides an answer, but does repeat content that is in the description.
@@
We used three types of evaluation methods for complementary insights on the quality of the model-generated answers. To obtain insights on the overall string similarity of the model-generated and the human-written answers, we use the well-established reference-based NLG metrics BLEU Papineni et al. (2002), CIDEr Vedantam et al. (2015), METEOR Banerjee and Lavie (2005), and ROUGE Lin (2004). However, these metrics do not adequately capture object hallucinations Rohrbach et al. (2018), which is why we supplement this analysis with a human-subject study specifically aimed at capturing hallucinated content. Finally, to measure how closely an answer’s visual details matches with the image and associated visual description, we also include an analysis using CLIPScore Hessel et al. (2021).
@@
For all models, we conducted two main experiments. The contextual condition simulated the situation human participants were in, where the models had complete access to the image, description and situational scenario while answering the question.
We include our prompt for the contextual condition in Appendix A.2. In addition, we tested a baseline condition where the models only had access to the image and question, in order to assess the model capabilities for incorporating contextual information. All evaluations were conducted with the greedy decoding method to ensure reproducibility.
@@
In Table 1, we show the performance of all models when provided with the image and question (baseline) and the full context that was available to human answerers (contextual). The IDEFICS model with the contextual condition has the highest performance for all models across all metrics. Evidently, integrating contextual information made the content itself more aligned with the ground-truth answer references, as evaluated by the reference-based metrics. While IDEFICS significantly improves when prompted with the contextual condition, this pattern is reversed for all other models. We hypothesize that less well-performing models closely reiterate the visual information they receive from the image description rather than adding new information to answer the question, which we turn to next.
5.1 Repetition of Visual Information Within Generated Answers
In order to investigate how much visual information models re-iterate within their generated answers, we conducted a similarity analysis of the model-generated answers to the image based on CLIPScore Hessel et al. (2021) and a similarity analysis of the generated answers to the human-written image descriptions using Sentence-BERT embeddings Reimers and Gurevych (2019). The analyses show convergent results.
While the reference-based metrics tend to decrease in the contextual condition, CLIPScore ratings (i.e., the similarity of the generated answer to the image) largely increase, as seen in Table 2. The only exception is BLIP-2, which significantly deteriorates in performance across metrics when contextual information is provided, largely rendering it uninterpretable.
Similarly, Figure 4 shows that the similarity between the description and the generated answers significantly increases with the contextual condition across models, according to a two-sample t-test analysis. However, the increase in description similarity between the baseline and contextual condition is the lowest for IDEFICS. We conclude that in their answers, the contextual versions of LLaVA, mPLUG-Owl, and BLIP-2 mimic the description more than IDEFICS does, which explains the CLIPScore increase across the baseline and contextual conditions for these three models. These findings suggest that those models might overly emphasize the descriptive details when available, leading to less alignment with the ground-truth answers.
@@ 5.2 Quantifying Hallucinations
While IDEFICS (contextual) best approximates the human answers, it’s still far from achieving human-like performance. When a model is intended to make information available that can’t be verified by a user, it is especially important that this model only generates truthful content MacLeod et al. (2017). While this is straightforward to evaluate for accuracy-based VQA datasets, this isn’t easily captured by the similarity-based metrics that are used for long-form evaluation, as measured by BLEU and CIDEr. In this section, we aim to give a brief intuition about the rate of hallucinated or incorrect pieces of information contained in the answers of the best-performing model, IDEFICS (contextual).
Model CLIPScore CLIPScore (C.)
IDEFICS 0.679 0.691
LLaVA 0.731 0.786
mPLUG-Owl 0.720 0.758
BLIP-2 0.613 0.586
Table 2: CLIPScore Improvement Across Baseline and Contextual Conditions. When each model is prompted with the contextual condition, CLIPScore increases, indicating that the model is repeating visual details. “C.” represents the contextual condition.
To estimate how many answers contain wrong, hallucinated or unverifiable content generated by the model, we conducted a human-subject study where participants were asked to rate model-generated answers. We randomly selected 100 answers from IDEFICS (contextual) to provide an assessment of the best-performing model. 70 participants were recruited, and compensated at a rate of \$13.80/hr. For each image, participants were asked to rate whether each answer contained information that was clearly not within the image, and to evaluate whether the image was strictly necessary for providing an answer to the question.
@@
Overall, participants indicated that 23% of the model-generated answers contained clearly erroneous information (with a Fleiss kappa inter-annotator agreement of 0.47) and for another 22%, the truthfulness of the answers couldn’t be clearly established. These results indicate that even the best-performing model (IDEFICS (contextual)) generates a high degree of erroneous information, making it unreliable for downstream use.
5.3 Evaluating Unanswerability
Since the questions in CommVQA are asked by people who cannot see the image, these questions are not always answerable. When a person asks an unanswerable question, ideally, models should abstain from answering rather than provide incorrect, hallucinated information Whitehead et al. (2022).
We evaluated all models on the 238 unanswerable questions in CommVQA. For each question, we assessed whether the model was able to successfully abstain from answering or if the model provided a hallucinated answer. To classify responses as “Abstention” or “Non-Abstention,” we conducted a direct string matching analysis to search for language that models commonly use to abstain (e.g., “I cannot answer.”) A full list of strings in this analysis is provided in Appendix B.
Figure 4: SBert Cosine Similarity Analysis in Human and Model Responses. Significance levels are marked with asterisks based on a two-sample t-test analysis.
Overall, IDEFICS (contextual) had the highest rate of successful abstentions at 21%, and BLIP-2 had the lowest rate, at 0% (in both contextual and baseline conditions). Without access to contextual information, IDEFICS’s abstention rate drops to 14%, suggesting that context is helpful for accurately identifying unanswerable questions. Appending the string “If you don’t know, say ‘unanswerable”’ to the model prompt improved the rate of IDEFICS’ (contextual) abstention performances on the unanswerable questions to 87%, suggesting that implicit instructions could help models abstain when necessary.
Questions most human participants judged as answerable tended to be answered by the models as well. Without the contextual condition, LLaVA and IDEFICS declined to answer for answerable questions at a rate of 8.3% and 8.6%, respectively. For both models, the rate of false positives decreased in the contextual condition (LLaVa: 7.6%; IDEFICS: 7.1%).
@@ 6 Conclusion
During the image description generation phase, initial descriptions are generated by GPT-4 and then edited by humans. This raises the potential issue that humans may be biased towards the GPT-4 responses, including keeping in details that may not be factually accurate. There is in fact evidence that starting out from an alt text shapes what type of descriptions people write and the quality of them Mack et al. (2021). Work on alt text captioning has primarily focused on old systems, which were far less competent, and more research is now needed to investigate how that interacts with newer systems. In our dataset, the descriptions themselves are not proposed as potentially ideal descriptions for accessibility. They simply provide the contextual framing for the questions after.
@@ Appendix A Appendix
A.1 Statistical Comparisons with Other VQA Datasets
In Table 3, we provide a statistical comparison of CommVQA with other VQA datasets.
Dataset Who Asked? $\overline{Q}$ $\overline{A}$ $\overline{C}$ Imgs # Scenarios # QA Pairs
VQA-v2 Crowdworkers 6.1 1.2 204,700 1 658,111
VizWiz-VQA Blind people 6.7 1.7 20,500 1 31,000
OK-VQA Crowdworkers 8.1 1.3 14,031 1 14,055
DocVQA Remote workers 9.5 2.4 1
Visual Dialog Crowdworkers 5.1 2.9 120,000 1 1.2M
VQAOnline Stack Overflow users 9.3 173.2 127 64,700 1 64,700
ScienceQA Elementary and high school curricula 12.1 4.4 41 6,500 1 21,208
CommVQA (ours) Crowdworkers who can only see the description 7.3 10.98 17.2 1000 6 8949
Table 3: Statistical Comparison: CommVQA and Other VQA Datasets. This table compares CommVQA with seven other VQA datasets. The first four rows cover image-question-answer inputs, while the bottom three rows cover image-question-answer-context inputs. $\overline{Q}$, $\overline{A}$, and $\overline{C}$ denote the average answer, question, and context lengths (as measured by the number of words), respectively. *Since Visual Dialog contains multi-turn conversations, we only included the average question length for the first round in the conversation.
A.2 Dataset Collection Overview
@@
For the scenario matching stage, we prompted GPT-4V with this prompt:
@@
For the answer elicitation study, we asked participants to: “Imagine you are browsing a scenario website when you encounter the following image. image Another user cannot see the image directly but has access to the following image description: description. Based on the description, they asked a follow-up question. Please answer based on the image: question.”
We recruited annotators through Prolific [3], an online annotation platform similar to Amazon’s Mechanical Turk. We recruited only US-based participants, and the only prescreener we used was to exclude participants who had taken a previous stage of the description generation process. For instance, participants who edited descriptions were excluded from providing questions or answers.
A.3 Context Integration Prompts for Models
@@
A.4 Additional Dataset Composition Analysis
Figure 5: CommVQA Dataset Examples. Four example entries from the CommVQA dataset, each paired with a randomly selected answer.
Figure 6: Examples of Model-Generated Answers. This figure shows examples of dataset entries and model responses from IDEFICS, the leading model. Sample responses from IDEFICS, the leading model, indicate that context aids in producing answers that better match human ground-truths, despite some limitations in context integration.
A.5 Models Chosen
We evaluated the following model versions, which are freely available on HuggingFace:
LLaVA
llava-v1.5-14b-3GB
IDEFICS
idefics-9b-instruct
mPLUG-Owl
mPLUG-owl-llama-7b
BLIP-2
blip2-opt-2.7b
Appendix B Unanswerability Analysis
In this section, we include a list of the strings used to evaluate whether a model abstained from answering.
  • I cannot answer
  • I cannot determine
  • I cannot provide
  • unanswerable
  • I don’t have enough information
  • AI language model
  • I don’t have the context
  • I don’t have enough context
  • I’m sorry
  • I don’t have the capability
  • I don’t have enough information
Model BLEU-1 BLEU-4 METEOR ROUGE CIDEr
IDEFICS 0.326 0.106 0.189 0.401 0.825
IDEFICS (Contextual) 0.356 0.122 0.198 0.413 0.904
mPLUG-Owl 0.191 0.045 0.166 0.311 0.414
mPLUG-Owl (Contextual) 0.178 0.043 0.169 0.284 0.335
LLaVA 0.192 0.050 0.171 0.311 0.431
LLaVA (Contextual) 0.149 0.037 0.161 0.247 0.200
BLIP-2 0.289 0.067 0.100 0.286 0.435
BLIP-2 (Contextual) 0.059 0.010 0.078 0.224 0.343
Table 4: Model Benchmarking Results on a Subset of 150 Images from CommVQA. Reflecting similar results from the full dataset benchmark, the highest-scoring model is IDEFICS (contextual), and LLaVA and mPLUG-Owl both exhibit decreased performance within the contextual condition.
Appendix C Effects of Dataset Size
We conducted an analysis on a subset of our dataset and uncovered similar effects of model behavior. In Table 4, we present model benchmarking results on a subset of 150 images and 1,215 unique visual questions within our dataset. Largely, we surface similar patterns as found within the full dataset. Namely, IDEFICS (contextual) exhibits the highest performance, and LLaVA and mPLUG-Owl both decrease their performance with their contextual condition, but this is explained by an increase in CLIPScore, as shown in Table 5.
Model CLIPScore CLIPScore (C.)
IDEFICS 0.667 0.666
LLaVA 0.729 0.784
mPLUG-Owl 0.720 0.756
BLIP-2 0.604 0.662
Table 5: CLIPScore Improvement Across Baseline and Contextual Conditions for Subset of 150 images. This table displays the increase in CLIPScore for each model when comparing baseline and contextual conditions for the subset of 150 images, finding a similar trend as when this result is run on the full dataset. In particular, LLaVA (C.) has the highest CLIPScore condition, and the CLIPScore stays stable across IDEFICS and IDEFICS (contextual).
Appendix D Potential Use Cases of the CommVQA Dataset
One potential use case for the CommVQA dataset is to assess whether a model performs better within a certain scenario than others. For instance, given a general-purpose VQA model, the CommVQA dataset provides a way to analyze whether this model excels at questions asked within the Shopping scenario, but struggles on questions asked within the Science Magazines scenario. These insights could prove useful for assessing when to deploy general-purpose VQA models versus specialized, domain-specific models.
Another use case is to evaluate a model’s ability to integrate contextual information along with the image. Prior work shows that people value visual explanations that incorporate contextual information Stangl et al. (2021, 2020); Muehlbradt and Kane (2022). But how well do models integrate this contextual information in practice? In Figure 4, we find that LlaVA and mPLUG-Owl both tend to repeat information from the description, while IDEFICS is more successful at integrating the contextual condition. CommVQA can help assess the ability of VQA models to integrate this contextual information, which as prior work has shown, is crucial for ensuring that model-generated outputs align with human preferences.
Appendix E Clarifying Context
This section serves to clarify the role of context and how it is distinct from prior work. In our case, “context” means not only the scenario (i.e., the website where the image was encountered), but the scenario and the description. Prior work suggests that for blind and low vision users, this broader context should shape people’s informational needs about an image Stangl et al. (2021, 2020). Although other types of context have been shown to result in an improvement in the VQA and image captioning literature, this pragmatic framing of context is unique to our dataset. Our work bridges the gap between the studies showing that this version of context matters, and puts this observation into practice in generating a VQA dataset with a more realistic distribution of questions and answers.